OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. y = x2 + 3x + 2 y = x2 + 2x + 1 Part 2: Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 3x + 2 different from y = x2 + 2x + 1?

hero (hero):

y = y x^2+3x + 2 = x^2 + 2x + 1 3x + 2 = 2x + 1 x = -1 y = 0 So they intersect at the point(-1,0)

hero (hero):

That's what they have in common

OpenStudy (anonymous):

Thankyou. Is that the answer?

hero (hero):

They also have different vertexes, different axes of symmetry and different x and y intercepts

hero (hero):

In addition to that, they have different factors. Similarities 1. Both quadratics 2. Both intersect at same point (-1,0) Differences: 1. Different axes of symmetry 2. Different x and y intercept 3. Different factors

OpenStudy (anonymous):

Okay thankss

hero (hero):

Well.... y = x2 + 3x + 2 factors to (x+1)(x+2) y = x2 + 2x + 1 factors to (x+1)(x+1) so they actually do have a common factor. That's another similiarity

OpenStudy (anonymous):

That's all I need but thankyou!

hero (hero):

Actually make sure you add that they have the common factor. That's a key thing to include.

OpenStudy (anonymous):

For which value of x does the graph of y = 2x2 − 7x + 6 cross the x-axis?

hero (hero):

Do you know how to factor that?

OpenStudy (anonymous):

yeah

hero (hero):

Well factor it....

hero (hero):

But post it on here

OpenStudy (anonymous):

7/4

hero (hero):

y = 2x2 − 7x + 6 y = 2x^2-4x-3x+6 y = 2x(x-2)-3(x-2) y=(x-2)(2x-3) x = 2, 3/2

OpenStudy (anonymous):

thanks. How can you tell when a quadratic equation has no real solutions?

hero (hero):

If the value of the discriminant is a negative number, then the quadratic has no real solutions